What is the length of a rectangle if the width is 10 cm and the diagonal is 16 cm?
Thank you!

Answer by checkley77(12844)

10^2+x^2=16^2
100+x^2=256
x^2=256-100
x^2=156
x=sqrt156
x=12.49 cm. ans. for the side.